OpenStudy (anonymous):

integral of cos5x cos3x dx

OpenStudy (anonymous):

Use the identity \(\cos x\cos y= \frac{1}{2}[\cos(x-y)+\cos(x+y)]\), we have \[\cos5x \cos3x = \frac{1}{2}[\cos(5x-3x)+\cos(5x+3x)]\] Then, integrate it. It shouldn't be difficult.

OpenStudy (mathmale):

Welcome to OpenStudy! Often trig identities can be used to simplify integration. In RolyPoly's second equation, combine: 5x-3x and 5x+3x. Re-write that second equation in simplified form and then ask yourself how you'd integrate the two resulting terms.